If Earthquake 1 measures 6.2 on the Richter scale, and Earthquake 2 measures 8.2 on the Richter scale, how many times bigger was Earthquake 2 than Earthquake 1? (Rephrased: By what factor was Earthquake 2 bigger than Earthquake 1?): If Earthquake 1 measures 6.2 on the Richter scale, and Earthquake 2 measures 8.2 on the Richter scale, how many times bigger was Earthquake 2 than Earthquake 1? (Rephrased: By what factor was Earthquake 2 bigger than Earthquake 1?)
Answer by Fombitz(1740)

The Richter scale is a logarithmic scale.
You subtract the two values.
The value is then raised to the power of ten.
x=8.2-6.2=2.0
D=10^x
D=10^(2.0)
D=100
A magnitude 8.2 earthquake is 100 times stronger than a magnitude 6.2 earthquake.